# MMMOS: Multi-domain Multi-axis Audio Quality Assessment

### v1

@@ Abstract
Index Terms:
Audio Quality assessment, Mean opinion score (MOS)
@@ I Introduction
This work presents our automatic quality assessment system MMMOS for AudioMOS Challenge 2025 track 2. This task aims to evaluate samples from text-to-speech (TTS), text-to-audio (TTA), and text-to-music (TTM) systems using 4 orthogonal axes to reduce the ambiguity of single-score evaluations. Production Quality (PQ) measures technical fidelity through clarity, dynamic range, frequency balance, and spatialization. Production Complexity (PC) quantifies the complexity of audio scenes by counting distinct sound components. Content Enjoyment (CE) captures the subjective aesthetic appeal in terms of emotional impact, artistic expression, and listener experience. Content Usefulness (CU) evaluates the suitability of samples as reusable material for content creation.
@@ II Related work
For general audio, a GRU-based approach [9] is first proposed to evaluate a single subjective quality score across speech, music, and environmental sounds. Meta’s Audiobox Aesthetics trains a multi-faceted assessment model of any audio clip across four dimensions using the WavLM [10] encoder, and calibrates the scores to be comparable for speech, music, and general audio. However, these works do not leverage the large-scale pretrained audio encoders from different domains.
Some recent studies aim to learn a single audio representation that spans multiple domains. Wu et al. [11] proposed concatenating latent features from speech and pitch encoders to create a “holistic” embedding. Ritter-Gutierrez et al. [12, 13] explore model-merging techniques to fuse speech and music encoders. While these methods improve general‐purpose audio representations, they have not been applied to perceptual quality assessment.
@@ III-A Dataset
We train and validate our models on two datasets: AES-natural [14], the official challenge dataset, and AES-PAM [15]. AES-natural aggregates audio from real-world, public corpora covering speech (EARS [16], LibriTTS [17], Common Voice 13.0 [18]), music (MUSDB18‐HQ [19], MusicCaps [20]), and sound effects (Audioset [21]). It covers different audio types, qualities, and sampling rates. AES-PAM, in turn, consists of synthetic utterances generated by various TTA and TTS systems, mirroring the artificial nature of the challenge test set. To preserve each method’s relative representation in both sets, we stratified the data by generation model and then partitioned each stratum into training and development subsets using an 80/20 split.
After downloading AES-natural’s clips, we exclude the audio clips unavailable on the web from our splits, yielding 3,367 training and 434 development samples. All audio is resampled to 16 kHz mono before feature extraction. During training, any clip longer than 10 seconds is randomly cropped to a 10-second segment, providing varied temporal contexts.
@@ III-B Model
Each waveform is passed in parallel through three frozen encoders: WavLM Base+ [^1] https://github.com/microsoft/unilm/tree/master/wavlm[10] for speech, MuQ [^2] https://huggingface.co/OpenMuQ/MuQ-large-msd-iter[22] for music, and M2D [^3] https://github.com/nttcslab/m2d?tab=readme-ov-file[23] for general audio. We extract the final hidden representations from each encoder at each time step and apply normalization to obtain fixed-length features. These normalized embeddings are concatenated along the feature axis to form a unified frame-level sequence.
@@ IV AudioMOS Challenge performance
Configuration Dev Set (utterance level) PAM Set (utterance level) Test Set (system level)
@@
MLP CCC 0.346 0.834 0.846 0.662 0.339 0.900 0.906 0.731 1.366 0.918 0.879 0.737
MLP CCC 0.399 0.814 0.830 0.652 0.284 0.907 0.913 0.741 1.146 0.938 0.912 0.779
BLSTM(t) CCC 0.351 0.833 0.842 0.660 0.277 0.906 0.906 0.732 1.307 0.937 0.901 0.760
BLSTM(h) CCC 0.406 0.810 0.815 0.638 0.301 0.903 0.909 0.739 1.360 0.919 0.888 0.746
MLP Con 0.390 0.825 0.839 0.663 0.329 0.912 0.915 0.746 1.679 0.930 0.890 0.733
MLP Con 0.415 0.813 0.820 0.645 0.329 0.907 0.912 0.739 1.576 0.939 0.916 0.789
BLSTM(t) Con 0.389 0.817 0.828 0.651 0.223 0.914 0.914 0.742 0.670 0.938 0.909 0.774
BLSTM(h) Con 0.347 0.827 0.838 0.668 0.229 0.917 0.916 0.750 0.986 0.939 0.912 0.777
MLP DCQ 0.421 0.817 0.832 0.650 0.329 0.907 0.911 0.741 1.617 0.921 0.887 0.744
MLP DCQ 0.448 0.818 0.824 0.645 0.279 0.918 0.921 0.754 1.466 0.944 0.919 0.796
BLSTM(t) DCQ 0.564 0.824 0.832 0.656 0.405 0.917 0.916 0.748 0.716 0.947 0.915 0.779
BLSTM(h) DCQ 0.515 0.822 0.831 0.650 0.427 0.903 0.900 0.724 0.974 0.939 0.909 0.773
MLP UT 0.361 0.822 0.849 0.672 0.298 0.910 0.911 0.745 1.163 0.919 0.872 0.726
MLP UT 0.342 0.833 0.838 0.656 0.268 0.909 0.912 0.744 1.112 0.940 0.913 0.782
BLSTM(t) UT 0.378 0.827 0.832 0.656 0.327 0.912 0.914 0.747 1.397 0.925 0.893 0.746
BLSTM(h) UT 0.358 0.823 0.836 0.653 0.245 0.916 0.919 0.752 1.113 0.945 0.920 0.793
@@ V-B Ablation on Adaptor
We use utterance‐level SRCC on the Dev set as our model‐selection criterion. As shown in Table III, the four downstream aggregation schemes differ by less than 0.02 in Dev set SRCC, and vary by under 0.01 in PAM dev set SRCC. Given these minimal differences, we include all four methods in subsequent experiments. On the unseen test set, the hidden‐state aggregation exhibits the strongest generalization, while the linear aggregator applied to BiEnc alone achieves only 0.872 system‐level SRCC, indicating comparatively poorer performance.
@@ V-C Ablation on Loss
Examining Table III, we observe that under the fixed encoder backbone (TriEnc) and aggregation scheme, the UTMOS loss achieves the highest utterance‐level SRCC on the Dev set (0.849) and consistently ranks first or second across all aggregation methods. However, on the unseen test set, UTMOS yields the top system‐level SRCC in only one aggregation configuration and falls behind DCQ and Con in the others, suggesting that its generalization performance is poor. In contrast, the DCQ and Con losses, although modest on the Dev set, generalize more robustly, attaining equal or better system‐level SRCC on the test set across all aggregators.
@@ VI Conclusion
We present a unified, non-intrusive audio quality assessment system that fuses features from three pretrained encoders (WavLM, MuQ, and M2D), evaluates three aggregation strategies and four loss functions, and ensembles top-performing models. In the AudioMOS 2025 Track 2 challenge, our system placed first on 6 of 8 production-complexity metrics and placed top-3 on 17 out of 32 metrics. Compared to the official baseline, we achieve a 20–30 % reduction in MSE and a 4–5 % increase in KTAU across all four perceptual axes.

### v2

@@ I Introduction
This work presents our automatic quality assessment system MMMOS [^1] https://github.com/robert0518/MMMOS for AudioMOS Challenge 2025 track 2. This task aims to evaluate samples from text-to-speech (TTS), text-to-audio (TTA), and text-to-music (TTM) systems using 4 orthogonal axes to reduce the ambiguity of single-score evaluations. Production Quality (PQ) measures technical fidelity through clarity, dynamic range, frequency balance, and spatialization. Production Complexity (PC) quantifies the complexity of audio scenes by counting distinct sound components. Content Enjoyment (CE) captures the subjective aesthetic appeal in terms of emotional impact, artistic expression, and listener experience. Content Usefulness (CU) evaluates the suitability of samples as reusable material for content creation.
@@ II Related work
For general audio, a GRU-based approach [13] is first proposed to evaluate a single subjective quality score across speech, music, and environmental sounds. Audiobox Aesthetics trains a multi-faceted assessment model of audio clips across four dimensions using the WavLM [4] encoder, and calibrates the scores to be comparable for speech, music, and general audio. However, these works don’t leverage the large-scale pretrained encoders from different domains.
Some recent studies aim to learn a single audio representation that spans multiple domains. Wu et al. [23] proposed concatenating latent features from speech and pitch encoders to create a “holistic” embedding. Ritter-Gutierrez et al. [18, 19] explore model-merging techniques to fuse speech and music encoders. While these methods improve general‐purpose audio representations, they havn’t been applied to perceptual quality assessment.
@@ III-A Dataset
We train and validate our models on two datasets: AES-natural [21], the official challenge dataset, and AES-PAM [5]. AES-natural aggregates audio from real-world, public corpora covering speech (EARS [17], LibriTTS [26], Common Voice 13.0 [2]), music (MUSDB18‐HQ [15], MusicCaps [1]), and sound effects (Audioset [7]). It covers different audio types, qualities, and sampling rates. AES-PAM, in turn, consists of synthetic utterances generated by various TTA and TTS systems, mirroring the artificial nature of the challenge test set. To preserve each method’s representation in both sets, we stratified the data by generation model and then partitioned each stratum into training and development sets using an 80/20 split.
After downloading AES-natural’s clips, we exclude the unavailable ones from our splits, yielding 3,367 training and 434 development samples. All audio is resampled to 16 kHz mono. During training, any clip longer than 10 seconds is randomly cropped to a 10-second segment, providing varied temporal contexts.
@@ III-B Model
Each waveform is passed in parallel through three frozen encoders: WavLM Base+ [^2] https://github.com/microsoft/unilm/tree/master/wavlm[4] for speech, MuQ [^3] https://huggingface.co/OpenMuQ/MuQ-large-msd-iter[28] for music, and M2D [^4] https://github.com/nttcslab/m2d/releases/download/v0.3.0/m2d_clap_vit_base-80x1001p16x16-240128_AS-FT_enconly.zip[14] for general audio. We obtain utterance-level embeddings from each encoder by performing mean pooling over the temporal dimension of their hidden representations. The resulting embeddings are then normalized individually to ensure comparable scales across encoders, and concatenated along the feature axis to form the final fused representation.
@@ IV AudioMOS Challenge performance
# Configuration Dev Set (utterance level) PAM Set (utterance level) Test Set (system level)
@@
1 MLP CCC 0.346 0.834 0.846 0.662 0.339 0.900 0.906 0.731 1.366 0.918 0.879 0.737
2 MLP CCC 0.399 0.814 0.830 0.652 0.284 0.907 0.913 0.741 1.146 0.938 0.912 0.779
3 BLSTM(t) CCC 0.351 0.833 0.842 0.660 0.277 0.906 0.906 0.732 1.307 0.937 0.901 0.760
4 BLSTM(h) CCC 0.406 0.810 0.815 0.638 0.301 0.903 0.909 0.739 1.360 0.919 0.888 0.746
5 MLP Con 0.390 0.825 0.839 0.663 0.329 0.912 0.915 0.746 1.679 0.930 0.890 0.733
6 MLP Con 0.415 0.813 0.820 0.645 0.329 0.907 0.912 0.739 1.576 0.939 0.916 0.789
7 BLSTM(t) Con 0.389 0.817 0.828 0.651 0.223 0.914 0.914 0.742 0.670 0.938 0.909 0.774
8 BLSTM(h) Con 0.347 0.827 0.838 0.668 0.229 0.917 0.916 0.750 0.986 0.939 0.912 0.777
9 MLP DCQ 0.421 0.817 0.832 0.650 0.329 0.907 0.911 0.741 1.617 0.921 0.887 0.744
10 MLP DCQ 0.448 0.818 0.824 0.645 0.279 0.918 0.921 0.754 1.466 0.944 0.919 0.796
11 BLSTM(t) DCQ 0.564 0.824 0.832 0.656 0.405 0.917 0.916 0.748 0.716 0.947 0.915 0.779
12 BLSTM(h) DCQ 0.515 0.822 0.831 0.650 0.427 0.903 0.900 0.724 0.974 0.939 0.909 0.773
13 MLP UT 0.361 0.822 0.849 0.672 0.298 0.910 0.911 0.745 1.163 0.919 0.872 0.726
14 MLP UT 0.342 0.833 0.838 0.656 0.268 0.909 0.912 0.744 1.112 0.940 0.913 0.782
15 BLSTM(t) UT 0.378 0.827 0.832 0.656 0.327 0.912 0.914 0.747 1.397 0.925 0.893 0.746
16 BLSTM(h) UT 0.358 0.823 0.836 0.653 0.245 0.916 0.919 0.752 1.113 0.945 0.920 0.793
@@ V-B Ablation on Adaptor
We use utterance‐level SRCC on the Dev set as our model‐selection criterion. As shown in Table III, the four downstream aggregation schemes (#13-16) differ by less than 0.02 in Dev set SRCC, and vary by under 0.01 in PAM dev set SRCC. Given these minimal differences, we include all four methods in subsequent experiments. On the unseen test set, the hidden‐state aggregation exhibits the strongest generalization, while the linear aggregator applied to BiEnc alone achieves only 0.872 system‐level SRCC, indicating comparatively poorer performance.
@@ V-C Ablation on Loss
Examining Table III, we observe that with the BiEnc backbone, the UTMOS loss achieves the highest utterance‐level SRCC on the Dev set (0.849). Under the TriEnc backbone, UTMOS loss consistently ranks first or second across all aggregation methods.However, on the unseen test set, UTMOS yields the top system‐level SRCC in only one aggregation configuration and falls behind DCQ and Con in the others, suggesting that its generalization performance is poor. In contrast, the DCQ and Con losses, although modest on the Dev set, generalize more robustly, attaining equal or better system‐level SRCC on the test set across all aggregators.
@@ VI Conclusion
We present a unified, non-intrusive audio quality assessment system that fuses features from three pretrained encoders, evaluates three aggregation strategies and four loss functions, and ensembles top-performing models. In the AudioMOS 2025 Track 2 challenge, our system placed first on 6 of 8 production-complexity metrics and placed top-3 on 17 out of 32 metrics. Compared to the official baseline, we achieve a 20–30 % reduction in MSE and a 4–5 % increase in KTAU across all four perceptual axes.